A box contains 4 red balls, 6 black balls, 8 blue balls and 12 white balls. What proportion of the balls are:
a) black?
b) red or blue?

Answer by macston(5194)

Total number of balls=30
a) proportion that are black balls =6/30=1/5
b) proportion that are red or blue =12/30=2/5